What is the probability that, for the sample of 115 customers, the proportion who wait longer than five days for their order to be delivered is less than 13.75%? Interpret your answer


1
Expert's answer
2022-05-23T23:35:37-0400

It has been found that 14% of customers of an online retail store wait longer than five days for their order to be delivered. We assume normal distribution


"\mu=0.14,"

"\sigma=\sqrt{\dfrac{p(1-p)}{n}}=\sqrt{\dfrac{0.14(1-0.14)}{115}}=0.0323567"

"P(X<0.1375)=P(Z<\dfrac{0.1375-0.14}{0.0323567})"

"\approx P(Z<-0.07726)\approx0.4692"

The probability that the proportion who wait longer than five days for their order to be delivered is less than 13.75% is 0.4692.